OpenStudy (anonymous):

what is the equation of the line in slope intercept form that passes through the point 3 -5 and is perpendicular to the line y = 3x - 6

OpenStudy (tcp2315):

perpendicular means it will have the negative reciprocal slope value, so m=-1/3 then just plug the point (3, -5) into the equation y=mx+b to solve for b and you're done so.. -5=(-1/3)3+b -5=(-1)+b -4=b therefore y=(-1/3)x-4